• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT
    Make New Entry, Make Followup Entry

    User name Meziani

    Log entry time 00:02:29 on February19,2011

    Entry number 347546

    keyword=Shift Summary

    Start of shift

    SL: Meziani, TO:Chao Gu, 3rd: B. Norum
    4:00 pm Moller measurement underway
    4:50 pm Beam charge asymmetry test
    6:05 pm Moller measurement completed
    7:00 MCC is sending CW into the Hall.
    8:20 Beam in the Hall at 5 uA in nominal position and raster size.
    Taking a run to check raster size: run 1528 on empty target
    Rebooted KODA
    8:40 pm Move target to BeO and check raster size
    run 1529
    8:47 pm Moving target to LD2 4cm (deuterium)
    Rates checked and OK for 20uA.
    9:00 pm MCC needs 15 minutes for ion chamber calibration.
    9:30 ion Chambers calibration completed.

    Start taking data
    run # 1531 20 uA
    Noticed quench of Q2 right arm around 10:00 pm
    Learned that the Quench happened in the previous shift before the Moller measurement. No one seems to have noticed and no information was passed on. The alarm handler was turn off too.

    run# 1531 was taken with Q2 down and Q3 with the wrong value of current 1176A

    10:20 Accessing the Hall to reset Q2 right arm.
    11:10 All set to take data at 20 uA.

    run # 1533 171K
    run # 1534 600K
    run # 1535

    12:00 pm shift Ends during run 1535